Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 213
Entry time: 07 Apr 2021, 08:52
|
Author: |
Andreas Galler |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 07.04.2021 |
|
|
General Announcement:
CTRL:
- recurrent error reprted by LAS --> drift compensation in SA2 --> need feedback
- minor issues reported instruments during startup
DA:
EEE:
- Issue on April 4th --> 4 PLCs shut down without a cause, (looks like power cut)
- MID crate went down at 9:00
- SA3 utca shut down (could be voltage drop, could be ehternet) (9:35)
VAC:
- VAC-OCD recovered SA2 vacuum state (valves, and ystem) (see EEE SA2 incident above)
- VAC obseved that the MID Mono warmed up, the cryo warmed up presumably after a PLCs update which interrupted an interloop communication
XPD:
DET:
ITDM:
- SA1 baclony room checked
- some racks lost power --> investifation going on (possibly happened on Sat. Apr 3rd 8:00 am)
XRO:
- incorrect motor velocity in SA3 --> fixed
-
LAS:
- small issue in SA2 with slow feedback --> see CTRL report
UNSYS:
- problem on SA1 central control node, after a power cycle aircoils and phase shifters were unresponsive
SPB/SFX:
- Laser Interlock issue was solved, power spike involved that reported by MEA3 ( klaus Witt).
FXE:
MID:
- M1 mirror was moved completely out on 1st of Apr.
- --> Naresh Kujala moved M1 out.
HED:
- Laser interlock issue on Tuesday
- with regards to the MID Mono cryo, there had been HED PLC updates
- calibration pipeline issue on AGIPD
- HED utca in optics hutch went off
SCS:
SQS:
|